1. Evaluate the integral using FTC. -1/2 (a) d. d. (b) 5°2+ (c) [(4-2)(1 – 4) dt (a) Lisa (cos (e) – e=%) dt
Answer #1

Amaner 9,3 The du use the common integral S store dx = (aresinews] 12 compute the boundaries limatot (aresin ()) Faresin (0)b> 2 tun du 4 expand and get tu 2 tu du 4 apply the sum rule 9 9 s 22 du + f F du 4 2 h du 4 = 2., h du 4 Commons integral Iput. n=4 = ln (4) = 2 10 (2) (by log reele) limuoa (losul) putreza Ilm (9) = 2 lm (3) (by log rule) » 211 (3) - 2 lm (2) = 2/and get simplify = [25] 4 compute the boundaries limaat + - (200) put us 4 2 54 limusa (en) put naq = 259 6 >> 6-4 now, from4 c.) (4-2) (4-4) at - expound and get H2-66f8 4 42-64 a 8 dd Apply sum/ difference nele 4 4 4 f t² dd - / 6A dd & J & dd -1compute the boundaries and get 15 2 = 6, 15 2 - 6. = 45 4 به هر 8 da 19 [st] 1) fostegral of a constant compute the boundarie12 d) (Cos (H)-e-ety at TA Apply sum rule 지고 cos(4) dt e-24 de TA I coa (A) dt TYA [sim (d)] / (by common postegral) T/A compadjust fortegral boundaries. > u=-7 T daa » usat = np noe = - 7 -14- Ludu -(- 4 [eu]-*12 if ) by common fostegral simplify an/, 2 J2 ел t e 41 2 et